UT 20170920 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (LBTO)
Observer AO-Team: SEsposito, EPinna, APuglisi, FRossi (Arcetri), DMiller, GTaylor (LBTO)
Telescope Operator: SAllenson

Observations done remotely from Tucson

Plan for the night

Same as last night, ie

- AO-teams working on optimization of FLAOs until ~3AM

- LUCI-AO takes over, short images checkout bino, perhaps incl. Bin=1

Summary

- Poor seeing all night, record was 4.96" on the DIMM (John needs to modify his LBTPlot smile )

- A tiny bit of AO-work

- No LUCI-AO

- JHill did some other tests

Details

All times are UT
Files are luci1/2.20170920.NNNN.fits

01:25 Sunset, wind close to the limit, about 15-18m/s, sky partly cloudy, hum = 75%, T=9.5deg

01:50 Low-level clouds have passed, Steve opens up, checks pointing and collimates

02:25 Gosh...the DIMM shows 2" seeing, the AO-guys try to use the bad seeing reconstructor

04:00 Sky is clear now, DIMM-seeing varies between 1" and 1.5", wind still around 15-18 m/s

04:30 DIMM-seeing stabilizes around 1"

06:15 DIMM-seeing varying wildly between 1" and 2.5" during the last 1.5hrs

08:30 The last two hours the DIMM-seeing varied wildly between 1" and 5" !!!!!

Handover to JHill for active optics / AGw tests relevant for PEPSI, no point in doing anything AO-relevant

11:21 DIMM-seeing has improved and stabilized around 1.2", we give it a try

11:33 Preset to SU Aurigae, bino-mode, Bin=1, OK, RunAO, failed,

11:58 Resent preset, RunAO failed again, seeing too unstable (mostly above 1.2"), stopped

-- %USERSIG{JochenHeidt - 2017-09-20}%
Topic revision: r6 - 20 Sep 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback